# Investigating Length Issues in Document-level Machine Translation

### v1

@@ A.5 Detailed Evaluation Results
sent FT. 85 86 85 85
Unif. 85 86 85 85
256 FT. 83 84 83 83
Unif. 83 84 83 82
512 FT. 82 84 83 82
Unif. 82 84 83 82
768 FT. 82 83 83 82
Unif. 82 84 83 82
1024 FT. 81 81 83 81
Unif. 81 83 83 81
1200 FT. 80 82 81 81
Unif. 80 82 81 81
1600 FT. 79 79 80 80
Unif. 80 80 78 79
2048 FT. 68 72 69 64
Unif. 70 72 69 67
Table 15: ds-BLEU (left) and 100×100$\times$100 ×COMET (right) for FT-Tower-U (FT. ) and Unif-Tower-U (Unif-full) trained on TED-U with target max source document length 2048 ($M=4096$).
TED-U TED-G FT Unif
FT vs Unif FT vs Unif U vs G U vs G
sent -0.1 (0.20) -0.0 (0.60) -0.1 (0.05) -0.1 (0.19)
256 0.5 (0.32) 0.1 (0.22) 0.5 (0.32) 0.1 (0.21)
512 0.3 (0.55) 0.1 (0.08) 0.1 (0.87) -0.1 (0.82)
768 0.2 (0.46) 0.0 (0.84) 0.3 (0.11) 0.2 (0.65)
1024 -0.2 (0.82) -0.4 (0.44) 0.6 (0.29) 0.4 (0.34)
1200 0.2 (0.44) 0.0 (0.70) 0.3 (0.24) 0.1 (0.71)
1600 0.6 (0.45) 0.4 (0.58) 0.1 (0.84) 0.0 (1.00)
2048 -0.5 (0.61) 0.2 (0.84) 0.7 (0.46) 1.3 (0.13)
TED-U TED-G FT Unif
FT vs Unif FT vs Unif U vs G U vs G
sent -0.0 (0.52) 0.0 (0.38) -0.0 (0.95) 0.0 (0.24)
256 0.0 (0.60) 0.0 (0.73) -0.0 (0.63) -0.1 (0.46)
512 -0.1 (0.15) 0.1 (0.16) -0.2 (0.09) 0.0 (0.76)
768 -0.1 (0.29) -0.1 (0.34) 0.0 (0.87) 0.0 (0.69)
1024 -0.6 (0.22) -0.3 (0.46) 0.1 (0.45) 0.3 (0.10)
1200 0.1 (0.69) 0.1 (0.50) 0.1 (0.61) 0.0 (0.93)
1600 0.2 (0.69) 0.6 (0.37) -0.2 (0.62) 0.2 (0.73)
2048 -0.9 (0.40) -0.2 (0.84) 0.9 (0.50) 1.5 (0.10)
Table 16: Average difference (and p-value of paired t-test) of ds-BLEU (top) evaluated on concatenated full TED talk and 100×100$\times$100 ×COMET (bottom) evaluated on realigned sentences, between FT-TowerBase (FT) and Unif-TowerBase (Unif) trained on TED-U (U) or TED-G (G).
Analysing the Repetitions
The analysis of $n$-gram repetitions is given in Table 9.

### v2

@@ A.5 Detailed Evaluation Results
sent FT 85 86 85 85
Unif 85 86 85 85
SHAPE 85 86 85 86
256 FT 83 84 83 83
Unif 83 84 83 82
SHAPE 83 84 83 82
512 FT 82 84 83 82
Unif 82 84 83 82
SHAPE 82 84 83 82
768 FT 82 83 83 82
Unif 82 84 83 82
SHAPE 82 83 83 82
1024 FT 81 81 83 81
Unif 81 83 83 81
SHAPE 81 82 83 81
1200 FT 80 82 81 81
Unif 80 82 81 81
SHAPE 80 82 81 80
1600 FT 79 79 80 80
Unif 80 80 78 79
SHAPE 79 81 78 80
2048 FT 68 72 69 64
Unif 70 72 69 67
SHAPE 70 73 70 68
Table 15: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) for FT-Tower-U (FT), Unif-Tower-U (Unif) and SHAPE-Tower-U (SHAPE) trained on TED-U with target max source document length 2048 ($M=4096$).
TED-U TED-G FT Unif SHAPE
FT vs Unif FT vs SHAPE FT vs Unif FT vs SHAPE U vs G U vs G U vs G
sent -0.1 (0.20) -0.2 (0.01) -0.0 (0.60) -0.1 (0.18) -0.1 (0.05) -0.1 (0.19) -0.1 (0.36)
256 0.5 (0.32) -0.0 (0.87) 0.1 (0.22) -0.5 (0.24) 0.5 (0.32) 0.1 (0.21) -0.1 (0.48)
512 0.3 (0.55) -0.1 (0.10) 0.1 (0.08) -0.6 (0.23) 0.1 (0.87) -0.1 (0.82) -0.4 (0.28)
768 0.2 (0.46) -0.2 (0.67) 0.0 (0.84) -0.9 (0.05) 0.3 (0.11) 0.2 (0.65) -0.4 (0.19)
1024 -0.2 (0.82) -0.3 (0.49) -0.4 (0.44) -0.5 (0.35) 0.6 (0.29) 0.4 (0.34) 0.5 (0.38)
1200 0.2 (0.44) 0.1 (0.82) 0.0 (0.70) -0.4 (0.06) 0.3 (0.24) 0.1 (0.71) -0.2 (0.50)
1600 0.6 (0.45) -0.7 (0.01) 0.4 (0.58) -0.5 (0.40) 0.1 (0.84) 0.0 (1.00) 0.4 (0.49)
2048 -0.5 (0.61) -1.0 (0.22) 0.2 (0.84) -1.3 (0.19) 0.7 (0.46) 1.3 (0.13) 0.4 (0.44)
sent -0.0 (0.52) -0.0 (0.93) 0.0 (0.38) 0.0 (0.60) -0.0 (0.95) 0.0 (0.24) 0.0 (0.56)
256 0.0 (0.60) 0.0 (0.85) 0.0 (0.73) -0.1 (0.07) -0.0 (0.63) -0.1 (0.46) -0.2 (0.05)
512 -0.1 (0.15) -0.2 (0.04) 0.1 (0.16) -0.1 (0.27) -0.2 (0.09) 0.0 (0.76) -0.0 (0.62)
768 -0.1 (0.29) 0.1 (0.28) -0.1 (0.34) -0.3 (0.00) 0.0 (0.87) 0.0 (0.69) -0.4 (0.00)
1024 -0.6 (0.22) -0.5 (0.32) -0.3 (0.46) -0.2 (0.09) 0.1 (0.45) 0.3 (0.10) 0.4 (0.42)
1200 0.1 (0.69) 0.2 (0.51) 0.1 (0.50) -0.4 (0.16) 0.1 (0.61) 0.0 (0.93) -0.4 (0.12)
1600 0.2 (0.69) -0.4 (0.27) 0.6 (0.37) 0.1 (0.70) -0.2 (0.62) 0.2 (0.73) 0.4 (0.49)
2048 -0.9 (0.40) -1.3 (0.16) -0.2 (0.84) -1.7 (0.13) 0.9 (0.50) 1.5 (0.10) 0.4 (0.47)
Table 16: Average difference (and p-values) in ds-BLEU (top) evaluated on full TED talks and 100×100$\times$100 ×COMET (bottom) evaluated on realigned sentences for TowerBase. Left and middle: paired comparison between the original fine-tuning (FT), unifPE and SHAPE on TED-U and TED-G respectively. Right: differences between fine-tuning on TED-U (U) and TED-G (G). A positive value indicates that in the comparison pair, the translation of the first item achieves higher scores than that of the second. Significant differences with p-values $<0.05$ are in bold.